LIFETIMES OF THE Ã STATES OF THE C_{3}-NE AND C_{3}-AR COMPLEXES

YI-JEN WANG, CHIAO-WEI CHEN, ANTHONY J. MERER, \underline{YEN-CHU HSU}, Institute of Atomic and Molecular Sciences, Academia Sinica, P. O. Box 23-166, Taipei 10617, Taiwan, R. O. C..

Under similar experimental conditions to our previous paper, the fluorescence lifetimes of the Ã states of C_{3}-Ne and C_{3}-Ar have been determined. The uncertainties in the measured lifetimes are estimated to be better than 10%. About 50 perpendicular bands and 11 parallel bands of the Ã − X̃ system of C_{3} were used to probe the fluorescence lifetimes of the Ne and Ar complexes. It is not possible to reduce the backing pressure while measuring lifetimes of the complexes; we therefore reduced the average pressure of the chamber to 1−2 × 10^{−5} torr by lowering the repetition rate. No sign of predissociation was found in either complex. Transitions to high vibrational levels of the Ã state of C_{3}, above 4500 cm^{−1}, give weak laser-induced fluorescence signals, so that lifetime measurements are not possible for some of these bands. The lifetimes of the levels of the Ne complex lying below 4300 cm^{−1} are generally shorter than 300 ns, except for some levels with C-C stretch excitation. The lifetimes of the Ne complex behave quite similarly to those of C_{3}. In contrast, the lifetimes of the Ar complex do not follow the vibrational dependence of the lifetimes of C_{3} and C_{3}-Ne. We speculate that the complicated lifetime pattern observed in the Ar complex can be attributed to interactions with nearby triplet levels.